OpenStudy (anonymous):

Help please Solve the equation. 0.25 r – 0.125 + 0.5 r = 0.5 + r

OpenStudy (ghazi):

-0.25r= 0.625 ..r=-2.5

OpenStudy (anonymous):

wait how did you get that?

OpenStudy (anonymous):

the 0.625 lost me

OpenStudy (ghazi):

0.75r-0.125=0.5+r (0.75-1)r= 0.5+0.125=0.625 -0.25r= 0.625 r=-2.5

OpenStudy (ghazi):

is it clear?

OpenStudy (ghazi):

@gungirl

OpenStudy (anonymous):

Yea thank you

OpenStudy (ghazi):

:)